# Reconstructing Quantum States and Expectations via Dynamical Tomography

### v1

@@ I Introduction
The tomography problem, i.e. state reconstruction from observable averages, is connected - and can be seen in fact as a generalization - of the so-called quantum marginal problem, where the available data are the state marginals with respect to a given multipartite and locality structure [10, 11, 12, 13, 14]. Dynamical tomography can be used in that setting as well, in general extending the set of states that are uniquely reconstructable given their marginals. A preliminary set of results in this sense has been presented in [7], where the notion of states that are Uniquely Dynamically Determined among All (UDDA) states has been introduced. Observability analysis in control theory is exploited to assess when all state are UDDA, and to provide a reconstruction protocol via linear inversion.
@@
  • We revisit the feasibility analysis when the dynamics is unitary, providing a lower bound for the minimal number of observables needed for DQST given Hamiltonian dynamics in multipartite systems, as well as showing that a generic unitary dynamics is sufficient to ensure DQST when the lower bounds are met. Furthermnore, we show that a class of dissipative Lindblad dynamics is able to ensure DQST from a single observable for networks of qubits. While it has been shown using different, more abstract methods that generic open dynamics suffice [6], our analysis exploits properties of the Pauli basis and shows that generic, purely dissipative generators are sufficient. The full proof, being a rather technical detour, is presented in Appendix A.3.
  • We propose a first solution to a problem that was left as open in the literature: the determination of a protocol to select the observables and times involved in the data-acquisition phase necessary for the state reconstruction. To this aim, in Section VI.2 we outline a heuristic procedure that addresses this problem and selects a minimum-cardinality set of observables, together with times (among the available ones) at which measurement should be performed on the system. The obtained sequence aims to maximize the signal-to-noise ratio of the measurements and the quality of the estimate by selecting observables that become the “most orthogonal” to the already selected ones.
@@ III.1 General Dynamics
$\mathbb{E}_{\rho}(\Phi_{t}[X_{i}])-\mathbb{E}_{\sigma}(\Phi_{t}[X_{i}])=\mathbb{E}_{{\rho}-{\sigma}}(\Phi_{t}(X_{i}))=0$
@@ III.2 Linear matrix representation via vectorization
The system isOBSERVABLEFeasibility ofDQST $\forall\!\rho_{0}\!\in\!\mathcal{D}(\mathcal{H})$Every state isUDDA
@@ IV.2 Observability for Continuous-Time Semigroup dynamics
Proof of the proposition, based essentially on Krylov subspaces, can be found in [9, 22].
@@
Let ${\rm vec}{(\mathcal{L}(X_{t}))}=L\,x_{i}$, where $L=L_{H}+L_{d}$ with
$L_{H}=i(I\otimes H-H^{\top}\!\otimes I),$ (12)
$L_{d}=\sum_{k}\!L_{k}^{\top}\!\otimes L_{k}^{\dagger}-\frac{1}{2}(I\otimes L_{k}^{\dagger}L_{k}+(L_{k}^{\dagger}L_{k})^{\top}\!\otimes I),$
@@
$O_{c}^{\dagger}=\begin{bmatrix}X&L^{\dagger}X&\dots&L^{{\dagger}(d^{2}-1)}X\end{bmatrix}$. (13)
@@ VI.1 Mean square error analysis
To evaluate the quality of the described DQST via liner regression protocol, it is possible to compute the Mean Squared Error (MSE) [28] between the actual (vectorized) state of the system ${r}_{0}$ and its estimate $\hat{{r}}_{0}$. The MSE in the described setting can be computed as
@@ VI.2 Selection of observables and measurement times
In particular, accordingly to the second point above, we can search among all the sets of evolved observables $\mathcal{R}$ the sets $\mathcal{R}^{*}$ solving the optimization problem
@@
The minimization of the functional in (22) is however impractical, as we need to be able to evaluate every evolved observable corresponding to $\mathcal{X}$ at all time instants $\mathcal{T}$ and to consider in $\mathcal{R}$ every possible combination of such observables. Similar problems emerge in trying to optimize standard tomographic protocol, see [29]. Furthermore, notice that while $\mathcal{X}$ is a set of finite cardinality, $\mathcal{T}$ may have infinite cardinality. In view of the difficulty of solving (22), we propose an heuristic iterative procedure for determining the observables to be considered in $\mathcal{R}$.
The identification of a set $\mathcal{R}$ or small cardinality is fundamental to reduce the number of experiment and the resources required for state reconstruction. The idea behind the procedure we propose below is therefore to maximize the information that can be acquired with a single measurement of an observable in $\mathcal{R},$ given that a number of previous measurements have already been chosen or performed. This allows for an iterative optimization procedure, which might be suboptimal in general.
@@
and $\mathcal{R}_{k}=\mathcal{R}_{k-1}\cup X^{k}$. These steps select the observable and the meaurement time that maximizes the amount of “new” information, that is, orthogonal with respect to the already available one. Although this does not lead, in general, to the optimal solution of the minimization problem described above, it is a viable proxy in many practical situations.
Finally, we stop the procedure when $k=d^{2}-1$. In most applications, it is not possible to let the system evolve for long time periods before performing measurements. It is therefore meaningful to restrict the search space for time instants to the set $\mathcal{T}_{\leq T}=\{t\in\mathcal{T}|t\leq T\}$.
@@ VII.2 Observability analysis
Dissipative Dynamics: We consider now a second possible dynamics for the system which encompasses dissipative terms. We choose a local Lindblad generator with noise operators
$L_{i}$ $=\eta_{i}\sigma_{i}^{+}$ $i\in\{1,\dots,4\}$
$L_{4+i}$ $=\eta_{i}\sigma_{i}^{-}$ $i\in\{1,\dots,4\}.$
The Hamiltonian part of the dynamics is the same as before, with all the coefficients set to 1. In contrast to the previous scenario, it is sufficient to consider $\eta_{i}=1\ \forall i$ and the system is observable. Therefore, the addition of noise terms is beneficial for observability, as expected in light of the results in Section IV.4.
@@ VII.3 Measurement time optimization
Since the considered system is continuous-time Markovian, we can run the algorithm described in Section VI to find the pairs of measurement operators in $\mathcal{X}$ and times in $\mathcal{T}$ identifying a set $\mathcal{R}$ of operators that can be measured on the system to solve DQST. The Algorithm is run only for the system in Scenario 2 as observability, is a requirement for the Algorithm. The Algorithm was implemented in Matlab, in particular the optimization of the functional in (23) was carried out using the fmincon matlab routine with initial time parameter equal to 1e61e-6, moreover the search space for the optimizer was constrained to the positive real line. In contrast, the search space for the optimization problem in (24) is a discrete set of relatively small cardinality, therefore the problem was solved in a combinatorial way by comparing the cost function for all evolved operators.
In Figure 3 are depicted the couples of measurement operators indexes and evolution times found by the Algorithm. We observe that all the measurement operators in the set $\mathcal{X}$ are chosen multiple times. Only the operator with index 1 is chosen a single time, this index however correspond to the identity operator which is a fixed point for the dynamics. We observe that collecting a single outcome of an experiment to perform DQST do not require waiting large time as the maximum found time is $t_{max}\approx 12$. It is interesting to compare this with the slowest decay time of the dissipative dynamics, that is $\tau\approx 1/|\lambda_{2}|$, where $\lambda_{2}$ is the non-zero eigenvalue of the Lindblad operator with largest real part. For the considered system $\lambda_{2}=-0.1308$ and $\tau\approx 7,64$. Hence, the largest time selected by the algorithm is under $2\tau.$
@@
0224466881010121205510101515ttOperator Index
Figure 3: Example 1 Scenario 2 Dissipative Dynamics: Pairs of times and measurement operators to perform DQST identified by the Algorithm described in section VI.
[…]
Figure 4: Squared error ($\varepsilon^{2}_{\rho}$) scaling in function of the number $N$ of instances of the system involved in computing the estimate $\hat{\rho}$ of the state $\rho$ of the system. We considered three states $\rho_{S},\rho_{GKS}$ and $\rho_{Gibbs}$ as described in the main text.
@@ VIII Example II – Coupled electron-nuclear systems
$\mathcal{X}$ EENN
Figure 5: Example 2: coupled electron (E) nuclear (N) system. The dotted line indicates the subsystem on which the meaurement operators act nontrivially $\mathcal{X}$.
@@
[…]
Figure 6: Squared error ($\varepsilon_{z}^{2}$) scaling in function of the number $N$ of instances of the system involved in computing the estimate $\hat{z}$ of the target observable $Z$. In computing the expectations we considered three states $\rho_{S},\rho_{GKS}$ and $\rho_{Gibbs}$ as described in the main text.
@@ IX Conclusions
In this work we lay the foundation for a systematic exploration of dynamical tomography, that is, a general method that allows for reconstructing states and expectations for a given quantum systems when limited observables are available, but we can exploit a precise knowledge of its dynamics. Leveraging control-theoretic observability analysis, we characterize systems for which any unknown state can be reconstructed from repeated experiments in the presence of general CPTP dynamics, and provide linear algebraic tests to check for such property for Markovian ones. In our framework, we also recall and extended existing results that show how unitarily evolving systems do not allow for full state reconstruction from a single observable, while open dynamics can. In presence of parametric dynamics, we show that if the system is observable for a choice of parameters, it must be so for almost all of them, allowing for randomized trials to test the feasibility of DQST. When the full state is not available, we characterize the subset of observable whose expectations can still be reconstructed from data. Algorithms to select measurements and evolution times in order to maximize the amount of new information and improve the quality of the estimate are also provided. The framework and the proposed methods are showcased in two physically-motivated examples: a chain of interacting qubits (spin 1/2), and a coupled electron-nucleus system that may serve as a simplified model for NV centers in diamonds. In order to further develop the analysis of dynamical tomography, some additional opportunity for investigations might address, among others: (i) a detailed error analysis and study of robustness with respect to uncertainty in the dynamical description; (ii) the development of improved algorithms for observable selection, in particular for locality constrained systems; and (iii) a in-depth comparison and connection of our approach with shadow tomography methods [25].
@@ Acknowledgements.
F.T. acknowledges funding from the European Union - NextGenerationEU, within the National Center for HPC, Big Data and Quantum Computing (Project No. CN00000013, CN 1,Spoke 10). F.T and T.G. are partially supported by the Italian Ministry of University and Research under the PRIN project “Extracting essential information and dynamics from complex networks”, grant no. 2022MBC2EZ.

### v2

@@ I Introduction
The tomography problem, i.e. state reconstruction from observable averages, is connected - and can be seen in fact as a generalization - of the so-called quantum marginal problem, where the available data are the state marginals with respect to a given multipartite and locality structure [10, 11, 12, 13, 14]. Dynamical tomography can be used in that setting as well, in general extending the set of states that are uniquely reconstructable given their marginals. A preliminary set of results in this sense has been presented in [7], where the notion of states that are Uniquely Dynamically Determined among All (UDDA) states has been introduced: these are states that can be uniquely reconstructed from a set of available measurements performed on identical preparations of such state, also allowing for arbitrary-time evolution with a known (closed or open) dynamical model. Observability analysis in control theory is exploited to assess when all states are UDDA, and to provide a reconstruction protocol via linear inversion.
@@
  • We revisit the feasibility analysis when the dynamics is unitary, providing a lower bound for the minimal number of observables needed for DQST given Hamiltonian dynamics in multipartite systems, as well as showing that a generic unitary dynamics is sufficient to ensure DQST when the lower bounds are met. Furthermore, we show that a class of dissipative Lindblad dynamics is able to ensure DQST from a single observable for networks of qubits. While it has been shown using different, more abstract methods that generic open dynamics suffice [6], our analysis exploits properties of the Pauli basis and shows that generic, purely dissipative generators are sufficient. The full proof, being a rather technical detour, is presented in Appendix A.3.
  • We propose a protocol to select the observables and times involved in the data-acquisition phase necessary for the state reconstruction. To this aim, in Section VI.2 we outline a heuristic procedure that addresses this problem and selects a minimum-cardinality set of observables, together with times (among the available ones) at which measurement should be performed on the system. The obtained sequence aims to maximize the signal-to-noise ratio of the measurements and the quality of the estimate by selecting observables that become the “most orthogonal” to the already selected ones. Albeit sub-optimal, the proposed procedure significantly reduces the computational burden by introducing iterative selection, rather than exploring all possible observable sequences, and represents the first attempt to address a problem that was left as open in the literature.
@@ III.1 General Dynamics
$\mathbb{E}_{\rho}(\Phi_{t}(X_{i}))-\mathbb{E}_{\sigma}(\Phi_{t}(X_{i}))=\mathbb{E}_{{\rho}-{\sigma}}(\Phi_{t}(X_{i}))=0$
@@ III.2 Linear matrix representation via vectorization
Remark 1.
It is worth noting that the class of non-Markovian dynamics we consider is very general, covering virtually any (determinsitic, average) open system evolution under the hypothesis of an initially factorized state (see e.g. the standard derivation of open system dynamics in [17]). The analysis extends non-trivially previous results: Not all dynamics can be satisfactorily approximated by Markovian ones. In particular, the reduced dynamics of a system of interest with an Hamiltonian coupling with a finite-dimensional “environment”, due to Poincare‘ recurrences, exhibits strong non-Markovian behaviour, that is revival of the initial conditions on sufficiently long time scales. In discrete time, an extreme example is provided by a pair of identical systems, evolving independently with CPTP maps, whose state is swapped at the end of every time step.
@@ IV.2 Observability for Continuous-Time Semigroup dynamics
Proof of the proposition, based essentially on Krylov subspaces, can be found in [9] and [22, Chapter 3].
@@
Let ${\rm vec}{(\mathcal{L}(X_{t}))}=\hat{L}\,x_{i}$, where $\hat{L}=\hat{L}_{H}+\hat{L}_{d}$ with
$\hat{L}_{H}=i(I\otimes H-H^{\top}\!\otimes I),$ (12)
$\hat{L}_{d}=\sum_{k}\!L_{k}^{\top}\!\otimes L_{k}^{\dagger}-\frac{1}{2}(I\otimes L_{k}^{\dagger}L_{k}+(L_{k}^{\dagger}L_{k})^{\top}\!\otimes I),$
@@
$O_{c}^{\dagger}=\begin{bmatrix}X&\hat{L}^{\dagger}X&\dots&\hat{L}^{{\dagger}(d^{2}-1)}X\end{bmatrix}$. (13)
@@ VI.1 Mean square error analysis
To evaluate the quality of the described DQST via linear regression protocol, it is possible to compute the Mean Squared Error (MSE) [28] between the actual (vectorized) state of the system ${r}_{0}$ and its estimate $\hat{{r}}_{0}$. The MSE in the described setting can be computed as
@@ VI.2 Selection of observables and measurement times
According to the second point above, in order to find the minimal MSE for the protocol, we can search among all the sets of evolved observables $\mathcal{R}$ the sets $\mathcal{R}^{*}$ solving the optimization problem
@@
However, minimization of the functional in (22) is impractical, as we need to be able to evaluate every evolved observable corresponding to $\mathcal{X}$ at all time instants $\mathcal{T}$ and to consider in $\mathcal{R}$ every possible combination of such observables. Similar problems arise in trying to optimize standard tomographic protocols; see [29]. Furthermore, notice that while $\mathcal{X}$ is a set of finite cardinality, $\mathcal{T}$ may have infinite cardinality. In view of the difficulty of solving (22), we propose an heuristic iterative procedure for determining the observables to be considered in $\mathcal{R}$.
The identification of a set $\mathcal{R}$ of small cardinality is fundamental to reduce the number of experiment and the resources required for state reconstruction. The idea behind the procedure is therefore to maximize the information that can be acquired with a single measurement of an observable in $\mathcal{R},$ given that a number of previous measurements have already been chosen or performed. This allows for an iterative optimization procedure, which might be suboptimal in general.
@@
and $\mathcal{R}_{k}=\mathcal{R}_{k-1}\cup X^{k}$. These steps select the observable and the measurement time that maximizes the amount of “new” information, that is, orthogonal with respect to the already available one.
Finally, we stop the procedure when $k=d^{2}-1$.
It is worth noting that, in most applications, it is not possible to let the system evolve for long time periods before performing measurements. It is therefore meaningful to restrict the search space for time instants to the set $\mathcal{T}_{\leq T}=\{t\in\mathcal{T}|t\leq T\},$ further reducing the resources required to compute the solution to the observable selection problem.
Although this procedure does not lead, in general, to the optimal solution of the minimization problem described above, it is a viable proxy in many practical situations. In Appendix B.1 we provide numerical evidence that our heuristic algorithm yields good approximations to the optimal solution, while a discussion of its numerical complexity is included in Appendix B.2.
@@ VII.2 Observability analysis
Dissipative Dynamics: We consider now a second possible dynamics for the system which encompasses a dissipative term. We choose a single, local Lindblad generator with noise operator
$L=4\sigma_{4}^{-}$
The Hamiltonian part of the dynamics is the same as before, with all the coefficients set to 1. In contrast to the previous scenario the system is observable. Therefore, the addition of noise terms is beneficial for observability, as expected in light of the results in Section IV.4.
@@ VII.3 Measurement time optimization
Since the considered system is continuous-time Markovian, we can run the algorithm AOT described in Section VI to find the pairs of measurement operators in $\mathcal{X}$ and times in $\mathcal{T}$ identifying a set $\mathcal{R}$ of operators that can be measured on the system to solve DQST. The AOT is run only for the system in Scenario 2 as observability, is a requirement for the AOT. The AOT was implemented in Matlab, in particular the optimization of the functional in (23) was carried out using the fmincon matlab routine with initial time parameter equal to 1e61e-6, moreover the search space for the optimizer was constrained to the positive real line. In contrast, the search space for the optimization problem in (24) is a discrete set of relatively small cardinality, therefore the problem was solved in a combinatorial way by comparing the cost function for all evolved operators.
In Figure 3 are depicted the couples of measurement operators indexes and evolution times found by the AOT. We observe that all the measurement operators in the set $\mathcal{X}$ are chosen multiple times. Only the operator with index 1 is chosen a single time, this index however correspond to the identity operator which is a fixed point for the dynamics. We observe that collecting a single outcome of an experiment to perform DQST do not require waiting large time as the maximum found time is $t_{max}\approx 12$. It is interesting to compare this with the slowest decay time of the dissipative dynamics, that is $\tau\approx 1/|Re{(\lambda_{2})}|$, where $\lambda_{2}$ is the non-zero eigenvalue of the Lindblad operator with largest real part. For the considered system $\lambda_{2}=-0.1308$ and $\tau\approx 7,64$. Hence, the largest time selected by the AOT is under $2\tau.$
@@
Figure 3: Example 1 Scenario 2 Dissipative Dynamics: Pairs of times and measurement operators to perform DQST identified by the AOT described in section VI [30].
Figure 4: Squared error ($\varepsilon^{2}_{\rho}$) scaling in function of the number $N$ of instances of the system involved in computing the estimate $\hat{\rho}$ of the state $\rho$ of the system. We considered three states $\rho_{S},\rho_{GKS}$ and $\rho_{Gibbs}$ as described in the main text [30].
@@ VIII Example II – Coupled electron-nuclear systems
Figure 5: Example 2: coupled electron (E) nuclear (N) system. The dotted line indicates the subsystem on which the measurement operators act nontrivially $\mathcal{X}$.
@@
Figure 6: Squared error ($\varepsilon_{z}^{2}$) scaling in function of the number $N$ of instances of the system involved in computing the estimate $\hat{z}$ of the target observable $Z$. In computing the expectations we considered three states $\rho_{S},\rho_{GKS}$ and $\rho_{Gibbs}$ as described in the main text [30].
@@ IX Conclusions
In this work we lay the foundation for a systematic exploration of dynamical tomography, that is, a general method that allows for reconstructing states and expectations for a given quantum systems when limited observables are available, but we can exploit a precise knowledge of its dynamics. Leveraging control-theoretic observability analysis, we characterize systems for which any unknown state can be reconstructed from repeated experiments in the presence of general CPTP dynamics, and provide linear algebraic tests to check for such property for Markovian ones. In our framework, we also recall and extended existing results that show how unitarily evolving systems do not allow for full state reconstruction from a single observable, while open dynamics can. In presence of parametric dynamics, we show that if the system is observable for a choice of parameters, it must be so for almost all of them, allowing for randomized trials to test the feasibility of DQST. When the full state is not available, we characterize the subset of observable whose expectations can still be reconstructed from data. A procedure (AOT) to select measurements and evolution times in order to maximize the amount of new information and improve the quality of the estimate is provided, which shows to significantly improve the quality of the estimate with respect to worst-case and randomized choices in the simulated scenarios. The framework and the proposed methods are showcased in two physically-motivated examples: a chain of interacting qubits (spin 1/2), and a coupled electron-nucleus system that may serve as a simplified model for NV centers in diamonds. In order to further develop the analysis of dynamical tomography, some additional opportunity for investigations might address, among others: (i) a detailed error analysis and study of robustness with respect to uncertainty in the dynamical description; (ii) the development of improved algorithms for observable selection, in particular for locality constrained systems; and (iii) a in-depth comparison and connection of our approach with shadow tomography methods [25].
@@ Acknowledgements.
F.T.acknowledges funding from the European Union - NextGenerationEU, within the National Center for HPC, Big Data and Quantum Computing (Project No. CN00000013, CN 1, Spoke 10) and European Union’s Horizon Europe research and innovation programme under the project “Quantum Secure Networks Partnership” (QSNP, grant agreement No 101114043). F.T and T.G. are partially supported by the Italian Ministry of University and Research under the PRIN project “Extracting essential information and dynamics from complex networks”, grant No. 2022MBC2EZ.
@@ A.3 Reconstructing N-qubit states from a single observable and purely dissipative dynamics
Appendix B On the Measurement Selection Algorithm
B.1 Comparison of selected measurements with the optimal sequence
The procedure for selecting the evolved observables proposed in Section VI.2 does not in general lead to the optimal solution of the minimization problem (22). However, we present next numerical evidence that the procedure helps to significantly reduce the value of the cost functional
$\mathcal{J}_{\mathcal{R}},={\rm tr}{[({O}_{\mathcal{R}}^{\dagger}{O}_{\mathcal{R}})^{-1}]}$
with respect to the worst possible one, and obtain relatively comparable values to the optimal solution.
In the following, we consider discretized systems with dissipative dynamics. In this setting, it is reasonable to assume that observables evaluated after sufficiently long evolution times carry negligible additional information about the initial state. Accordingly, we restrict the search over evolution times to the finite interval $[0,2T]$, where $T$ denotes the slowest decay time of the dynamics, defined as
$T=\frac{1}{|Re(\lambda_{2})|},$
with $\lambda_{2}$ the non-zero eigenvalue of the Lindblad operator having the largest real part (the spectral gap). In all examples, we discretize this interval and take
$\mathcal{T}=\left\{\frac{2T\,k}{15}\;:\;k\in\{0,\dots,15\}\right\}.$
With this choice, it is numerically feasible to compute all evolved observables $X[t]$ with $X\in\mathcal{X}$ and $t\in\mathcal{T}$, yielding a finite pool of cardinality $q=|\mathcal{X}|\,|\mathcal{T}|$. Since DQST requires at least $d^{2}$ evolved observables, we can then evaluate the cost functional $J_{\mathcal{R}}$ over all $\binom{q}{d^{2}}$ possible selections of $d^{2}$ observables from this pool and determine its minimum.
Example 1: single qubit. We consider a qubit with Hilbert space $\mathcal{H}=\mathbb{C}^{2}$, and a Lindblad generator with Hamiltonian and single noise operator
$H=\sigma_{x},\qquad L=\sigma_{-},$
where $\sigma_{x}$ is the Pauli $x$ operator and $\sigma_{-}$ is the lowering operator. We assume the available measurement set
[…]
where $\left|0\right>=[1\ \ 0]^{\dagger}$ and $\left|+\right>=\frac{1}{\sqrt{2}}([1\ \ 0]^{\dagger}+[0\ \ 1]^{\dagger})$. The identity need not be included since the dynamics is trace preserving.
We used Matlab to build the sampled observability matrix and verify the Kalman rank condition (observability is required for DQST and hence for the applicability of the selection procedure). We then: (i) ran our heuristic selection to obtain a set $\mathcal{R}$ and its cost $J_{\mathcal{R}}^{A}$, (ii) computed the global minimum and maximum costs, $J_{\mathcal{R},{\rm min}}^{C}$ and $J_{\mathcal{R},{\rm max}}^{C}$, via exhaustive search over all $d^{2}=4$-subsets from the $q=32$ evolved observables, and (iii) computed the average randomized cost $J_{\mathcal{R}}^{S}$ over the retained randomized realizations.
The resulting values are:
$J_{\mathcal{R},{\rm min}}^{C}$ $=7.6\times 10^{-1},$ $J_{\mathcal{R}}^{A}$ $=1.4\times 10^{1},$
$J_{\mathcal{R},{\rm max}}^{C}$ $=1.0\times 10^{8},$ $J_{\mathcal{R}}^{S}$ $=7.2\times 10^{4}.$
These results confirm that the heuristic is sub-optimal (here $J_{\mathcal{R}}^{A}$ is about two orders of magnitude above $J_{\mathcal{R},{\rm min}}^{C}$), but it substantially improves over the worst-case selection and over the randomized benchmark. In particular, the relative gap with respect to the worst-case: $\frac{J_{\mathcal{R}}^{A}-J_{\mathcal{R},{\rm min}}^{C}}{J_{\mathcal{R},{max}}^{C}}=1.3\times 10^{-7},$ is very small, indicating that the algorithm yields a cost much closer to the best case than to the worst case on this instance. Moreover, $J_{\mathcal{R}}^{A}\ll J_{\mathcal{R}}^{S}$, showing that the heuristic improves (on average) over the simple randomized iterative procedure.
Example 2: two qubits. We consider two qubits, $\mathcal{H}=\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ with $\mathcal{H}_{q}=\mathbb{C}^{2}$ and $\mathcal{H}\simeq\mathbb{C}^{4}$. The considered Lindblad generator is associated to an Hamiltonian:
$H=\sum_{i=1}^{2}\big(\sigma_{i}^{x}+\sigma_{i}^{y}+\sigma_{i}^{z}\big)\;+\;\sigma_{1}^{x}\sigma_{2}^{x}+\sigma_{1}^{z}\sigma_{2}^{z},$ (35)
and noise operators
$L_{i}$ $=\frac{1}{\sqrt{2}}\sigma_{i}^{+},$ $L_{2+i}$ $=\frac{1}{\sqrt{2}}\sigma_{i}^{-},$ $i\in\{1,2\}.$
Here $\sigma_{j}^{\alpha}$ denotes the Pauli operator $\sigma^{\alpha}$, $\alpha\in\{0,x,y,z\}$, acting on the $j$-th qubit (e.g. $\sigma_{1}^{\alpha}=\sigma^{\alpha}\otimes I$, $\sigma_{2}^{\alpha}=I\otimes\sigma^{\alpha}$), and $\sigma_{j}^{\pm}$ are the raising/lowering operators on the $j$-th qubit.
We assume we can measure only the first qubit, with
$\mathcal{X}=\{\left|0\right>\left<0\right|\otimes I,\;\left|+\right>\left<+\right|\otimes I\}.$
As in Example 1, we verified observability and computed: the heuristic cost $J_{\mathcal{R}}^{A}$; and the exhaustive-search minimum/maximum costs $J_{\mathcal{R},{\rm min}}^{C}$, $J_{\mathcal{R},{\rm max}}^{C}$ over all $d^{2}=16$-subsets from the $q=32$ evolved observables that yield a full-rank observability matrix. The obtained values are:
$J_{\mathcal{R},{\rm min}}^{C}$ $=7.8\times 10^{5},$ $J_{\mathcal{R}}^{A}$ $=3.7\times 10^{6},$
$J_{\mathcal{R},{\rm max}}^{C}$ $=8.6\times 10^{18},$ $J_{\mathcal{R}}^{S}$ $=1.1\times 10^{11}$
The interpretation is analogous to Example 1: the heuristic is sub-optimal (here $J_{\mathcal{R}}^{A}$ is one order of magnitude above $J_{\mathcal{R},{\rm min}}^{C}$), but it yields a dramatic improvement compared with the worst-case selection and averaged cost. The relative gap with respect to the worst-case: $\frac{J_{\mathcal{R}}^{A}-J_{\mathcal{R},{\rm min}}^{C}}{J_{\mathcal{R},{max}}^{C}}=3.4\times 10^{-13},$ is again very small.
Overall, these results support the claim that the heuristic algorithm—while less demanding than exhaustive search—yields good approximations to the optimal solution in the tested instances.
B.2 Computational considerations
The procedure we propose is iterative and has the following peculiarities:
  1. 1.
    It involves $n^{2}$ steps, where $n=\dim(\mathcal{H})$.
  2. 2.
    In the first step one observable at random is chosen, this operation has cost $O(1)$
  3. 3.
    At every step k, with $k>1$, it requires to search the observable which is the most orthogonal to the one selected at the previous time step. More formally, let $\mathcal{R}_{k}$ be the set of measurement operators found in steps $1,\dots,k$ and $X^{k}\in\mathcal{X}$ the observable that is chosen in the $k$-th step of the procedure. In the $k$-th step we select the observable $X^{k}=X_{i}^{*}[t^{*}]$, with $X_{i}^{*}\in\mathcal{X}$, $t^{*}\in\mathcal{T}$, with maximum projection on $\textrm{span}\{\mathcal{R}_{k-1}\}^{\perp}$.
    Let $\Pi_{\mathcal{R}^{\perp}}$ be the orthogonal projector onto $\textrm{span}\{\mathcal{R}_{k-1}\}^{\perp}$, for all $X_{i}\in\mathcal{X}$ we first find
[…]
    where $\left|\left|\cdot\right|\right|_{HS}$ is the Hilbert-Schmidt norm.
    The cost of this operation is dominated by the computational cost of the algorithm chosen to perform the optimization of the functional: call this cost $A$. In practice the vectorized representation of the system can be used, therefore the evaluation of the functional requires only the multiplication of matrices (in $\mathbb{C}^{n^{2}\times n^{2}}$) and vectors (in $\mathbb{C}^{n^{2}}$) and matrix exponentiation (to compute the evolved observable ant time $t$). These are all operations polynomial in $n^{2}$ and at most O($n^{6}$). We emphasize that the optimization is performed for all the $|\mathcal{X}|$ observables in $\mathcal{X}$.
    Successively we set
[…]
    and $\mathcal{R}_{k}=\mathcal{R}_{k-1}\cup X^{k}$. This optimization requires selecting among the $|\mathcal{X}|$ observables output of the previous time step the one that minimizes the cost functional. The value of the cost functional for each of these observables has already been computed in the previous optimization step and can be stored in memory. Since no other evaluation of the functional is needed the computational cost of the operation is $O(|\mathcal{X}|)$.
Each of the $k$ step has therefore cost O($|\mathcal{X}|An^{6}$) the overall cost of the procedure is O($|\mathcal{X}|An^{8}$).